# FitzHugh-Nagumo equation: from global dynamics to the slow-fast system

### v1

@@ 4.1. The critical manifold and bifurcations
$A=\dfrac{-\dfrac{\partial l_{1}}{\partial x}+3\dfrac{\partial l_{2}}{\partial x}-2\dfrac{\partial l_{4}}{\partial x}+2l_{6}}{8}\qquad B=\dfrac{\dfrac{\partial l_{3}}{\partial x}+l_{6}}{2},$
where the functions $l_{i},\,i=1,\ldots,6$ and their partial derivatives are evaluated at the point $\left(x,y,\lambda,\varepsilon\right)=\left(0,0,0,0\right)$. The Hopf bifurcation is nondegenerate when $A\neq 0$, it is supercritical if $A<0$ and subcritical if $A>0$.
@@ 4.2. Case A $(a=0)$
The curves $T_{F}^{2,3}$, $T_{H}^{2,3}$, $DH$ and SNL, introduced in Theorem 3.4 and Remark 3.5, can be parameterised by $c\in[1,+\infty)$. From now on, we restrict our attention to the portions of these curves lying within the region defined by $b,c\geq 1$ – see Figure 6. We are going to denote by $T_{F,+}^{2,3}\backslash\{(1,1)\}$ the connected component of $T_{F}^{2,3}\backslash\{(1,1)\}$ with highest values of $b$ for the same $c$, and $T_{F,-}^{2,3}\backslash\{(1,1)\}$ the other. It is straightforward to verify that, for sufficiently large values of $c$ in $[1,+\infty)$, the following holds:
$T_{F,-}^{2,3}(c)\leq T_{H}^{2,3}(c)\leq DH(c)\leq SNL(c)\leq T_{F,+}^{2,3}.$ (4.8)
All the curves admit $(b,c)=(1,1)$ as accumulation point. Denote by $A_{1}\leq A_{2}<A_{3}<A_{4}$ the limits when $c\to+\infty$ of the previous curves (when it exists), respectively. More specifically, we have:
Figure 6. Partial scheme of the codimension 1 bifurcation curves emerging from the point $(b,c)=(1,1)$ associated with $f_{(0,b,c)}$ of (3.1), for $b,c>0$. The symbol $c^{\infty}$ represents the limit of the associated curves when $c\to+\infty$ (implicit compactification in $c$). The notation of the curves correspond to those of Theorem 3.4, for Case A. The line SNL is dashed because it does not make part of the DZ bifurcation with $\mathbb{Z}_{2}(\kappa)$–symmetry. Near the curve $b=T_{H}^{2,3}$, the emerging periodic solutions give rise to canards as a consequence of Proposition 4.9.
@@
With respect to (4.4), there exists a smooth function $b_{c}:\left[0,\varepsilon_{0}\right]\to\mathbb{R}$ that associates each value of $\varepsilon\in\left(0,\varepsilon_{0}\right]$ to a value $b$ that gives rise to a family of canards in the system, asymptotically defined by:
@@
The difference $b-3/2$ of system (4.4) is played by $\lambda$ of Theorem 4.7. The equilibrium $E_{3}$ is a regular singular fold point and there is a Hopf bifurcation point at $b=3/2$. In order to apply Theorem 4.7 we change variables and the parameter by
$\left\{\begin{array}[]{l}\bar{x}=x-1\\ \\ \bar{y}=y+2/3\\ \\ \lambda=b-3/2\\ \end{array}\right.$
@@
$\left\{\begin{array}[]{l}\dot{\bar{x}}=\bar{y}-\frac{1}{3}\bar{x}^{3}-\bar{x}^{2}\\ \\ \dot{\bar{y}}=\varepsilon\left(-\bar{x}-\lambda\bar{y}+\frac{2}{3}\lambda-\frac{3}{2}\bar{y}\right).\end{array}\right.$
In the notation of Theorem 4.7 we have
@@
hence $A=-1/2<0$ and $B=-3/4<0$ therefore the Hopf bifurcation is supercritical for $b$ in the original system (4.9). Since for $b<T_{H}^{2,3}(c)$ the eigenvalues of $Df_{(0,b,c)},$ evaluated at $E_{i},$ $i=1,2,$ are negative, then the bifurcating periodic solution is unstable and canards occur when
@@ 4.3. Digestive remark
We describe the dynamics of (3.1) in the case $a=0$, focusing on the first quadrant of the bifurcation diagram in the $(b,c)$–plane. We suggest the reader follows the description by using Figure 6.
For $b\in(0,1)$ and $c>1$, there exists a unique equilibrium point $E_{1}$, which is unstable. This source is enclosed by a stable periodic orbit denoted by $C^{s}$. At $b=1$, one of the eigenvalues of the source $E_{1}$ becomes zero, leading to the creation of two additional unstable nodes $E_{2}$ and $E_{3}$. $E_{1}$ undergoes a supercritical Pitchfork bifurcation.
The saddles $E_{2}$ and $E_{3}$ subsequently become unstable foci and undergo a supercritical Hopf bifurcation, giving rise to two unstable periodic solutions (symmetric under $\kappa$). Around $b=DH(c)$, the invariant manifolds of $E_{1}$ evolve into a double homoclinic orbit, which is unstable, previously identified in [8].
As $b$ increases smoothly, the double homoclinic loop breaks apart, resulting in an unstable periodic orbit $C^{u}$, which surrounds both the stable and unstable manifolds of $E_{1}$. Near $b=SNL(c)$, the periodic orbits $C^{s}$ and $C^{u}$ collapse, and two stable equilibria remain, dominating the basin of attraction of any compact subset of $\mathbb{R}^{2}$. For $c\gg 1$, there is a small portion of the bifurcation parameter diagram $(b,c)$ of Figure 6, where the periodic orbits emerging from $T_{H}^{2,3}$ correspond to canards. They rapidly (with respect to $b$) collapse into the double homoclinic cycle to $E_{1}$. The values of $b_{c}$ that give rise to canards are at most $\mathcal{O}\left(\varepsilon\right)$ away from the values $b_{H}$ where a Hopf bifurcation occurs (see Figure 7):
@@
Figure 7. Sketch of the singular bifurcations for $b\geq 1$ and $c=c^{\infty}$, in Case A. The smaller $\varepsilon$ is, the narrower the interval of $[b_{c},b_{H}]$. When $\varepsilon=0$, we have $b_{c}\equiv b_{H}$.
@@ 4.4. Case B $(b=0)$
With respect to (4.4), there exists a continuous function $a_{c}:\left[0,\varepsilon_{0}\right]\to\mathbb{R}$ that associates each value of $\varepsilon\in\left[0,\varepsilon_{0}\right]$ to a value $a$ that gives rise to a canard in the system, asymptotically defined by:
@@
The difference $a-1$ of system (4.4) is played by $\lambda$ in Theorem 4.7. We are interested in only one of the two fold points of (3.1) located at $(1,-2/3)$. After shifting the fold point to the origin, reversing time $t\to-t$ and setting $\lambda=a-1$, we get
@@ 4.5. Digestive remark
The dynamics of system (3.1) for $b=0$ are illustrated in Figures 4 and 9. Our analysis focuses on the first and fourth quadrants of the bifurcation diagram in the $(a,c)$–plane. For $a<-1$, the saddle $E_{1}$ becomes a stable node. Near $a=-1$, a Belyakov transition occurs, followed by a supercritical Hopf bifurcation that gives rise to a stable periodic orbit that involves into a canard in an exponentially small portion of the bifurcation plane. This periodic solution disappears through a subcritical Hopf bifurcation at $a=1$.
5. The phase portraits in the Poincaré disc
In this section, we study the global phase portraits for system (3.1) in the Poincaré disc by using the theory described in Subsections 2.2 and 2.3.
@@ 5. The phase portraits in the Poincaré disc
The global phase portraits of (3.1) in the Poincaré disc are topologically equivalent to one of the phase portraits of Figures 8, 9 and 10 for parameters values in Cases A, B and C, respectively.
Remark 5.2.
The calculations used in the proof of Theorem 5.1 could be streamlined and simplified by using weighted or quasi-homogeneous blow-ups. By a pedagogical point of view, we decided to do three blow-ups, which are illustrated in Figures 8, 9 and 10 for parameters values in Cases A, B and C. Some computations in the following proof were carried out using the software Maple.
@@
System (3.1) in the local chart $U_{1}=\{z_{1}>0\}$ may be written as:
@@
$\dot{v}=$ $-cuv^{3}-cv^{3}+\frac{c}{3}v.$
The origin is the unique equilibrium on $v=0$ (the points of $\mathbb{S}^{1}$ in any chart have $v=0.$) The eigenvalues of Jacobian matrix associated with the vector field (5.1) are $\dfrac{c}{3},\dfrac{c}{3}.$ So, the origin is an unstable node (resp. stable) for $c>0$ (resp. $c<0$.)
System (3.1) in the local chart $U_{2}$ may be written as:

### v3

@@ 4.1. The critical manifold and bifurcations
$A=-\dfrac{\dfrac{\partial l_{1}}{\partial x}-3\dfrac{\partial l_{2}}{\partial x}+2\dfrac{\partial l_{4}}{\partial x}-2l_{6}}{8},\qquad B=\dfrac{\dfrac{\partial l_{3}}{\partial x}+l_{6}}{2},$
and the functions $l_{i},$ for $1\leq i\leq 6$, and their partial derivatives are calculated at $\left(x,y,\lambda,\varepsilon\right)=\left(0,0,0,0\right)$. The latter bifurcation is non-degenerate when $A\neq 0$, it occurs for positive values of $\lambda-\lambda_{H}$ if $A<0$, and for negative values of $\lambda-\lambda_{H}$ if $A>0$.
@@ 4.2. Case A $(a=0)$
[…]
$T_{F,-}^{2,3}(c)\leq T_{H}^{2,3}(c)\leq DH(c)\leq SNL(c).$ (4.8)
All the curves pass through $(b,c)=(1,1)$. Denote by $A_{1}\leq A_{2}<A_{3}<A_{4}$ the limits when $c\to+\infty$ of the previous curves (when it exists), respectively. More specifically, we have:
Figure 6. Partial scheme of the codimension 1 bifurcation curves emerging from the point $(b,c)=(1,1)$ associated with $f_{(0,b,c)}$ of (3.1), for $b,c>0$. The symbol $c^{\infty}$ represents “very large” values of $c$. The notation of the curves correspond to those of Theorem 3.4, for Case A. The line SNL is dashed because it does not make part of the DZ bifurcation with $\mathbb{Z}_{2}(\kappa)$–symmetry. Near the curve $b=T_{H}^{2,3}$, the emerging periodic solutions give rise to canards as a consequence of Proposition 4.9.
@@
With respect to (4.4), a smooth function $b_{c}:\left[0,\varepsilon_{0}\right]\to\mathbb{R}$ exists that assigns every $\varepsilon\in\left(0,\varepsilon_{0}\right]$ to a value $b$ leading to a family of canards in the system, asymptotically defined by:
@@
The difference b3/2b-3/2 of system (4.4) is played by $\lambda$ of Theorem 4.7. The equilibrium $E_{3}$ is a singular fold point and a Hopf bifurcation occurs at $b=3/2$. Using Theorem 4.7, we perform a change of variables and parameter as
[…]
@@
$\left\{\begin{array}[]{l}\dot{\bar{x}}=\bar{y}-\bar{x}^{2}-\frac{1}{3}\bar{x}^{3}\\ \\ \dot{\bar{y}}=\varepsilon\left(-\bar{x}+\frac{2}{3}\lambda-\lambda\bar{y}-\frac{3}{2}\bar{y}\right).\end{array}\right$.
Using the notation introduced in Theorem 4.7, we obtain
@@
hence A=1/2<0A=-1/2<0 and B=3/4<0B=-3/4<0 therefore the Hopf bifurcation occurs for values $\lambda>\lambda_{H}(\Leftrightarrow b>\frac{3}{2}+\lambda_{H})$ . Since for $b<T_{H}^{2,3}(c)$ the eigenvalues of $Df_{(0,b,c)},$ evaluated at $E_{i},$ $i=1,2,$ are negative, then the bifurcating periodic solution is unstable and canards occur when
@@ 4.3. Digestive remark
We describe the dynamics of (3.1) in the case $a=0$, focusing on the first quadrant of the bifurcation diagram in the (b,c)(b,c)–plane. We suggest the reader follows the description by using Figure 6.
For $b\in(0,1)$ and $c>1$, there exists a unique equilibrium point $E_{1}$, which is unstable. This source is enclosed by a stable periodic solution denoted by $C^{s}$. At $b=1$, one of the eigenvalues of $Df_{(0,b,c)}$ at $E_{1}$ becomes zero, leading to the creation of two additional unstable nodes $E_{2}$ and $E_{3}$. In other words, $E_{1}$ undergoes a supercritical (with respect to $b$) Pitchfork bifurcation.
The saddles $E_{2}$ and $E_{3}$ subsequently become unstable foci and undergo a supercritical Hopf bifurcation, leading to two unstable periodic solutions (symmetric under $\kappa$). Around $b=DH(c)$, the invariant manifolds of $E_{1}$ evolve into a double homoclinic orbit, which is unstable, previously identified in [8].
As $b$ increases smoothly, the double homoclinic loop breaks apart, resulting in an unstable periodic orbit $C^{u}$, which surrounds both the stable and unstable manifolds of $E_{1}$. Near $b=SNL(c)$, the periodic orbits $C^{s}$ and $C^{u}$ collapse, and two stable equilibria remain, dominating the basin of attraction of any compact subset of $\mathbb{R}^{2}$. For $c\gg 1$, there is a small portion of the bifurcation parameter diagram $(b,c)$ of Figure 6, where the periodic orbits emerging from $T_{H}^{2,3}$ correspond to canards. They rapidly collapse into the double homoclinic cycle to $E_{1}$. Parameters $b_{c}$ that yield canards are only $\mathcal{O}\left(\varepsilon\right)$ away from the corresponding Hopf bifurcation values $b_{H}$ (see Figure 7):
@@
Figure 7. Sketch of the singular bifurcations for $b\geq 1$ and $c<\infty$, in Case A. As $\varepsilon$ decreases, the interval $b_{c},b_{H}]$ becomes narrower. When $\varepsilon=0\,(\Leftrightarrow c=+\infty$), we have $b_{c}\equiv b_{H}$.
@@ 4.4. Case B $(b=0)$
With respect to (4.4), a continuous mapping $a_{c}:\left[0,\varepsilon_{0}\right]\to\mathbb{R}$ exists that assigns to each $\varepsilon\in\left[0,\varepsilon_{0}\right]$ a corresponding value $a$ giving rise to a canard, defined asymptotically by
@@
The difference a1a-1 of system (4.4) is played by $\lambda$ in Theorem 4.7. We are interested in only one of the two fold points of (3.1) located at (1,2/3)(1,-2/3). After shifting the fold point to the origin, reversing time $t\to-t$ and setting $\lambda=a-1$, we get
@@ 4.5. Digestive remark
The dynamics of system (3.1) for $b=0$ is illustrated in Figures 4 and Table 2. Our analysis focuses on the first and fourth quadrants of the bifurcation diagram in the (a,c)(a,c)–plane and $c>0$. For a<1a<-1, the saddle $E_{1}$ becomes a stable node. Near a=1a=-1, a Belyakov transition occurs, followed by a supercritical Hopf bifurcation that gives rise to a stable periodic orbit that involves into a canard in an exponentially small portion of the bifurcation plane. This periodic solution disappears through another Hopf bifurcation at $a=1$.
5. The phase portraits “near infinity”
Studying the dynamics “near infinity” provides insight into the system’s asymptotic behaviour and allows the identification of trajectories that leave bounded sets. Using the techniques of Subsections 2.2 and 2.3, we compactify system (3.1) and describe the dynamics of (3.1) at infinity – this is what we call the dynamics “near infinity”. We distinguish between four cases
$b\geq 0,\,c>0,\qquad b\geq 0,\,c<0,\qquad b<0,\,c>0,\qquad b<0,\,c>0.$ (5.1)
The main theorem of this section is the following:
@@ 5. The phase portraits “near infinity”
The dynamics of system (3.1) “near infinity” is topologically equivalent to one of the phase portraits depicted in Figure 8.
Insight into the reasoning
Before proceeding with the formal proof of the result, we give a geometrical interpretation of what is going to be done: we investigate the behaviour of trajectories of the planar vector field (3.1) as their position tends toward infinity. To do this, we apply the Poincaré compactification of Subsection 2.3, which transforms points at infinity into a bounded region represented by a disk containing a finite number of equilibria. We then study the qualitative dynamics of the “compactified” vector field paying particular attention to the equilibria located there, the directions along which trajectories diverge, and the conditions under which the flow at infinity becomes attracting or repelling. Some equilibria are singularities. In order to study their dynamics we need to desingularize them using the techniques of Subsection 2.2.
(a) $b\geq 0$ and $c>0$
(b) $b\geq 0$ and $c<0$
(c) $b<0$ and $c<0$
(d) $b<0$ and $c>0$
Figure 8. Dynamics near infinity on the Poincaré disc of system (3.1).
@@
Using (2.3) with $d=3$, system (3.1) in the local chart
$U_{1}=\{(y_{1},y_{2},y_{3})=(x_{1},x_{2},1)\in\mathbb{S}^{2}:y_{1}>0\}$
may be written as:
@@
$\dot{v}=$ $-cuv^{3}-cv^{3}+\frac{c}{3}v,$
where
$u=\frac{x_{2}}{x_{1}}\quad\text{and}\quad v=\frac{1}{x_{1}}.$
The system (5.2) has a unique equilibrium at the origin on $v=0$ (in any chart, points of $\mathbb{S}^{1}$ are characterized by $v=0$). The eigenvalues of the Jacobian matrix associated to (5.2) computed at the origin, are $\dfrac{c}{3},\dfrac{c}{3}.$ So, the origin is an unstable (resp. stable) node for $c>0$ (resp. $c<0$).
Analogously, using (2.4) and $d=3$, the system (3.1) in chart $U_{2}$ takes the form: